OpenStudy (anonymous):

what are the first five terms of the sequence given by the formula an=4n=1

OpenStudy (anonymous):

ok so with sequences you will take an=4n+1 i'm assuming?

OpenStudy (anonymous):

i assume so

OpenStudy (anonymous):

well is it an=4n+1? or an=4n-1?

OpenStudy (anonymous):

+1

OpenStudy (anonymous):

ok so this is pretty easy so you will go an=4(1)+1=5 then an=4(2)+1=9 then an=4(3)+1= an=4(4)+1= an=4n(5)+1= you understand?

OpenStudy (anonymous):

im pretty sure i am

OpenStudy (anonymous):

ok awesome have a good day and make sure you give me best response ;)

OpenStudy (anonymous):

of course

OpenStudy (anonymous):

so................whats the answer